# Congestion-Approximators from the Bottom Up

### v2

@@ 2.1 Bottom-Up Construction
As mentioned previously, the partition $\mathcal{P}_{2}$ is a weak-expander decomposition of the graph, so it can be computed in nearly-linear time using off-the-shelf expander decomposition algorithm that avoid any black-box call to max-flow [34]. For partitions $\mathcal{P}_{3}$ onwards, expansion is measured with respect to a subset of edges, so simple expander decomposition algorithms no longer suffice. Instead, we have to resort to expander decomposition algorithms that make black-box calls to (approximate) max-flow. Naïvely, these max-flows can be computed recursively, resulting in a congestion approximator algorithm that makes recursive calls to max-flow, similar to [33]. Our key insight is that these max-flow instances are actually well-structured enough that recursion is unnecessary. In particular, to construct partition $\mathcal{P}_{i+1}$, the first $i$ partitions $\mathcal{P}_{1},\mathcal{P}_{2},\ldots,\mathcal{P}_{i}$—which the algorithm has already computed—can be converted to a pseudo-congestion approximator, which then translates to a max-flow algorithm sufficient for these well-structured instances.
@@ 3 Preliminaries
We are given an undirected, capacitated graph $G=(V,E)$. The graph has $n$ vertices and $m$ edges, and each edge $e\in E$ has capacity $c(e)$ in the range $[1,W]$. For a set $C\subseteq V$, let $\partial C$ denote the set of edges with exactly one endpoint in $C$, and let $\delta C=\sum_{e\in\partial C}c(e)$ denote the total capacity of edges in $\partial C$. For a vertex $v\in V$, let $\deg(v)$ denote the capacitated degree of vertex $v$, which is also equal to $\delta\{v\}$. When the graph $G$ is clear from context, we may drop the subscript $G$. We sometimes write $\partial_{G}C$, $\delta_{G}C$, and $\deg_{G}(v)$ to emphasize that the values are with respect to graph $G$.
@@
Given a graph $H=(U,F)$, a congestion-approximator $\mathcal{C}$ of quality $\alpha$ is a collection of subsets of $U$ such that for any demand $\mathbf{b}$ satisfying $|\mathbf{b}(C)|\leq\delta_{H}C$ for all $C\in\mathcal{C}$, there is a flow routing demand $\mathbf{b}$ with congestion $\alpha$. Through Sherman’s framework [36, 37], a congestion-approximator of quality $\alpha$ translates to a $(1+\epsilon)$-approximate maximum flow algorithm with running time $\tilde{O}(\epsilon^{-1}\alpha m)$. [^12] The $\tilde{O}(\cdot)$ notation hides polylogarithmic factors in $n$. In our paper, it is most convenient to work with a stronger variant of approximate min-cut/max-flow called fair cut/flow [25], which is formally defined in Section 5.2.
@@ 4 Bottom-Up Congestion-Approximator
In this section, we formally state our congestion-approximator construction and prove its quality guarantee. See Figure 1 for an illustration. The most important distinction is that we define and analyze mixing on vertex weightings, not edges. This simplifies the notation since we can avoid working with the subdivision graph in [33]. For example, in property (2) below, the mixing of the vertex weighting $\mathrm{deg}_{\partial\mathcal{P}_{i}\cup\partial C}|_{C}$ is conceptually equivalent to the mixing of the edges of $\partial\mathcal{P}_{i}$ internal to $C$ together with the boundary edges $\partial C$.
@@
Then, their union $\mathcal{C}=\bigcup_{i\in[L]}\mathcal{R}_{\geq i}$ is a congestion-approximator with quality $5L^{2}\alpha\beta$.
@@
Consider their union $\mathcal{C}=\bigcup_{i\in[L]}\mathcal{R}_{\geq i}$. Then, for any demand $\mathbf{b}\in\mathbb{R}^{V}$ satisfying $|\mathbf{b}(C)|\leq\delta C$ for all $C\in\mathcal{C}$, there exists a demand $\mathbf{b}^{\prime}\in\mathbb{R}^{V}$ satisfying $|\mathbf{b}^{\prime}|\leq L\deg_{\partial\mathcal{P}_{L}}$ and a flow routing demand $\mathbf{b}-\mathbf{b}^{\prime}$ with congestion $5L^{2}\alpha\beta$.
@@
Consider partitions $\mathcal{P}_{1},\mathcal{P}_{2},\ldots,\mathcal{P}_{L}$ that satisfy the assumptions of Theorem 4.1. For a given demand $\mathbf{b}\in\mathbb{R}^{V}$ satisfying $|\mathbf{b}(C)|\leq\delta C$ for all $C\in\mathcal{C}$, we want to establish a flow routing demand $\mathbf{b}$ with congestion $5L^{2}\alpha\beta$. Theorem 4.1 then follows from the definition of congestion-approximator.
Apply Lemma 4.2 to the partitions $\mathcal{P}_{1},\mathcal{P}_{2},\ldots,\mathcal{P}_{L}$ and the demand $\mathbf{b}$. We obtain a demand $\mathbf{b}^{\prime}\in\mathbb{R}^{V}$ satisfying $|\mathbf{b}^{\prime}|\leq L\deg_{\partial\mathcal{P}_{L}}$ and a flow $f$ routing demand $\mathbf{b}-\mathbf{b}^{\prime}$ with congestion $5L^{2}\alpha\beta$. By assumption (1) of Theorem 4.1, we have $\mathcal{P}_{L}=\{V\}$, which implies that $\partial\mathcal{P}_{L}=\emptyset$. Since $|\mathbf{b}^{\prime}|\leq L\deg_{\partial\mathcal{P}_{L}}=\mathbf{0}$, we must have $\mathbf{b}^{\prime}=\mathbf{0}$. It follows that the flow $f$ routes demand $\mathbf{b}$ with congestion $5L^{2}\alpha\beta$, finishing the proof. ∎
@@
Let $\mathbf{b}\in\mathbb{R}^{V}$ be a flow demand satisfying $|\mathbf{b}(C)|\leq\delta C$ for all $C\in\mathcal{C}$. We need to construct a demand $\mathbf{b}^{\prime}\in\mathbb{R}^{V}$ satisfying $|\mathbf{b}^{\prime}|\leq L\deg_{\partial\mathcal{P}_{L}}$ and a flow routing demand $\mathbf{b}-\mathbf{b}^{\prime}$ with congestion $5L^{2}\alpha\beta$.
@@
    The flow $f_{i}$ has congestion $5L\alpha\beta$.
@@
    The demand $\mathbf{b}_{i}$ satisfies $|\mathbf{b}_{i}|\leq(i+1)\deg_{\partial\mathcal{R}_{\geq i+1}}$.
Note for the last level $L-1$, by definition we have $\partial\mathcal{R}_{\geq L}=\partial\mathcal{P}_{L}$. The lemma below shows that properties (1), (2), and (4) alone are sufficient to prove Lemma 4.2 with demand $\mathbf{b}^{\prime}=\mathbf{b}_{L-1}$ and flow $f_{1}+f_{2}+\cdots+f_{L-1}$.
@@
Suppose that properties (1), (2), and (4) hold for each $i\in[L-1]$. Then, the demand $\mathbf{b}_{L-1}$ satisfies $|\mathbf{b}_{L-1}|\leq L\deg_{\partial\mathcal{P}_{L}}$, and the flow $f_{1}+f_{2}+\cdots+f_{L-1}$ routes demand $\mathbf{b}-\mathbf{b}_{L-1}$ with congestion $5L^{2}\alpha\beta$.
@@
The demand $\mathbf{b}_{L-1}$ satisfies $|\mathbf{b}_{L-1}|\leq L\deg_{\partial\mathcal{R}_{\geq L}}=L\deg_{\partial\mathcal{P}_{L}}$ by property (4) on iteration $i=L-1$. By property (1) over all $i\in[L-1]$, the flow $f_{1}+f_{2}+\cdots+f_{L-1}$ routes demand $(\mathbf{b}_{0}-\mathbf{b}_{1})+(\mathbf{b}_{1}-\mathbf{b}_{2})+\cdots+(\mathbf{b}_{L-2}-\mathbf{b}_{L-1})=\mathbf{b}-\mathbf{b}_{L-1}$. The congestion is at most the sum of the congestions of each $f_{i}$, which is $(L-1)\cdot 5L\alpha\beta\leq 5L^{2}\alpha\beta$ by property (2). ∎
@@
    $|\mathbf{s}|\leq i\deg_{\partial\mathcal{P}_{i}}$.
@@
    The flow $f$ has congestion $5L\alpha\beta$.
@@
For the inductive step, we apply Lemma 4.6 on iteration $i\geq 1$ and the vector $\mathbf{s}\in\mathbb{R}^{V}$ with
$\mathbf{s}(v)=\begin{cases}\displaystyle\left(1-\frac{\deg_{\partial\mathcal{R}_{\geq i+1}}(v)}{\deg_{\partial\mathcal{R}_{\geq i}}(v)}\right)\mathbf{b}_{i-1}(v)&\text{if }\deg_{\partial\mathcal{R}_{\geq i}}(v)>0,\\ 0&\text{otherwise}.\end{cases}.$
We first verify the conditions on $\mathbf{s}$ required by Lemma 4.6.
@@
    To establish condition a, fix a vertex $v\in V$; our goal is to show that $|\mathbf{s}(v)|\leq i\deg_{\partial\mathcal{P}_{i}}(v)$. If $\deg_{\partial\mathcal{R}_{\geq i}}(v)=0$, then $\mathbf{s}(v)=0$ and the claim is trivial. Otherwise, suppose that $\deg_{\partial\mathcal{R}_{\geq i}}(v)>0$. We can apply property (4) for iteration $i-1$ to obtain
    $|\mathbf{s}(v)|$ $=\left(1-\frac{\deg_{\partial\mathcal{R}_{\geq i+1}}(v)}{\deg_{\partial\mathcal{R}_{\geq i}}(v)}\right)|\mathbf{b}_{i-1}(v)|$
    $\leq\left(1-\frac{\deg_{\partial\mathcal{R}_{\geq i+1}}(v)}{\deg_{\partial\mathcal{R}_{\geq i}}(v)}\right)\cdot i\deg_{\partial\mathcal{R}_{\geq i}}(v)$
    $=(\deg_{\partial\mathcal{R}_{\geq i}}(v)-\deg_{\partial\mathcal{R}_{\geq i+1}}(v))\cdot i.$
    which equals $i\deg_{\partial\mathcal{R}_{\geq i}\setminus\partial\mathcal{R}_{\geq i+1}}(v)$ since $\partial\mathcal{R}_{\geq i}\supseteq\partial\mathcal{R}_{\geq i+1}$ by Claim 4.3. Since $\partial\mathcal{R}_{\geq i}\setminus\partial\mathcal{R}_{\geq i+1}\subseteq\partial\mathcal{P}_{i}$ by Claim 4.4, we have $\deg_{\partial\mathcal{R}_{\geq i}\setminus\partial\mathcal{R}_{\geq i+1}}(v)\leq\deg_{\partial\mathcal{P}_{i}}(v)$, establishing condition a.
@@
    By definition of vector $\mathbf{s}$, we have $|\mathbf{s}(C)|\leq|\mathbf{b}_{i-1}(C)|$, which equals $|\mathbf{b}_{0}(C)|$ from above. Finally, since the initial flow demand $\mathbf{b}\in\mathbb{R}^{V}$ satisfies $|\mathbf{b}(C)|\leq\delta C$, we have $|\mathbf{s}(C)|\leq|\mathbf{b}_{0}(C)|=|\mathbf{b}(C)|\leq\delta C$, establishing condition b.
With the conditions fulfilled, Lemma 4.6 outputs a flow $f$ which we set as $f_{i}$, immediately satisfying property (2). We set $\mathbf{b}_{i}=\mathbf{b}_{i-1}-\mathbf{s}+\mathbf{t}$ so that flow $f_{i}$ routes demand $\mathbf{s}-\mathbf{t}=\mathbf{b}_{i-1}-\mathbf{b}_{i}$ and property (1) holds. Property (3) follows from property iii of Lemma 4.6.
It remains to establish property (4). By property (4) for iteration $i-1$, we have $|\mathbf{b}_{i-1}(v)|\leq i\deg_{\partial\mathcal{R}_{\geq i}}(v)$ for each vertex $v\in V$. If $\deg_{\partial\mathcal{R}_{\geq i}}(v)=0$, then $\mathbf{b}_{i-1}(v)=0$ and
$|\mathbf{b}_{i}(v)|=|\mathbf{b}_{i-1}(v)-\mathbf{s}(v)+\mathbf{t}(v)|=|0-0+\mathbf{t}(v)|=|\mathbf{t}(v)|.$
Otherwise, if $\deg_{\partial\mathcal{R}_{\geq i}}(v)>0$, then
$|\mathbf{b}_{i}(v)|$ $=|\mathbf{b}_{i-1}(v)-\mathbf{s}(v)+\mathbf{t}(v)|$
$=\left|\frac{\deg_{\partial\mathcal{R}_{\geq i+1}}(v)}{\deg_{\partial\mathcal{R}_{\geq i}}(v)}\mathbf{b}_{i-1}(v)+\mathbf{t}(v)\right|$
$\leq\frac{\deg_{\partial\mathcal{R}_{\geq i+1}}(v)}{\deg_{\partial\mathcal{R}_{\geq i}}(v)}|\mathbf{b}_{i-1}(v)|+|\mathbf{t}(v)|$
$\leq\frac{\deg_{\partial\mathcal{R}_{\geq i+1}}(v)}{\deg_{\partial\mathcal{R}_{\geq i}}(v)}\cdot i\deg_{\partial\mathcal{R}_{\geq i}}(v)+|\mathbf{t}(v)|$
$=i\cdot\deg_{\partial\mathcal{R}_{\geq i+1}}(v)+|\mathbf{t}(v)|.$
In both cases, we obtain $|\mathbf{b}_{i}(v)|\leq i\cdot\deg_{\partial\mathcal{R}_{\geq i+1}}(v)+|\mathbf{t}(v)|$. We also have $|\mathbf{t}(v)|\leq\deg_{\partial\mathcal{R}_{\geq i+1}}(v)$ by property i of Lemma 4.6, so property (4) holds on iteration $i$, completing the induction. ∎
For the rest of this section, we establish Lemma 4.6, the most technical part of the proof. Before we do so, we first establish three technical helper claims.
@@
Claim 4.9.
For any $i\in[L-1]$, consider any vector $\mathbf{x}\in\mathbb{R}^{V}_{\geq 0}$ with $\mathbf{x}\leq\deg_{\partial\mathcal{R}_{\geq i+1}}$. There exists a vector $\mathbf{y}\in\mathbb{R}^{V}_{\geq 0}$ with $\mathbf{y}\leq\deg_{\partial\mathcal{P}_{i}}$ and a flow routing demand $\mathbf{x}-\mathbf{y}$ with congestion $(2L-1-2i)\beta$.
@@
We prove the statement by induction from $i=L-1$ down to $i=1$. For the base case $i=L-1$, since $\mathcal{R}_{\geq L}=\mathcal{P}_{L}$, we have $\mathbf{x}\leq\deg_{\partial\mathcal{P}_{L}}$. By Claim 4.8 on vector $\mathbf{x}$, there is a vector $\mathbf{y}\in\mathbb{R}^{V}_{\geq 0}$ with $\mathbf{y}\leq\deg_{\partial\mathcal{P}_{L-1}}/2\leq\deg_{\partial\mathcal{P}_{L-1}}$ and a flow routing demand $\mathbf{x}-\mathbf{y}$ with congestion $\beta=(2L-1-2(L-1))\beta$.
@@
For the inductive step, consider a vector $\mathbf{x}\in\mathbb{R}^{V}_{\geq 0}$ with $\mathbf{x}\leq\deg_{\partial\mathcal{R}_{\geq i+1}}$. Define $\mathbf{x}^{\prime}\in\mathbb{R}^{V}_{\geq 0}$ as
@@
$\mathbf{x}^{\prime}(v)=\begin{cases}\displaystyle\frac{\deg_{\partial\mathcal{R}_{\geq i+2}}(v)}{\deg_{\partial\mathcal{R}_{\geq i+1}}(v)}\cdot\mathbf{x}(v)&\text{if }\deg_{\partial\mathcal{R}_{\geq i+1}}(v)>0,\\ 0&\text{otherwise},\end{cases}$

### v3

@@ 2.1 Bottom-Up Construction
As mentioned previously, the partition $\mathcal{P}_{2}$ is a weak-expander decomposition of the graph, so it can be computed in nearly-linear time using off-the-shelf expander decomposition algorithm that avoid any black-box call to max-flow [SW19]. For partitions $\mathcal{P}_{3}$ onwards, expansion is measured with respect to a subset of edges, so simple expander decomposition algorithms no longer suffice. Instead, we have to resort to expander decomposition algorithms that make black-box calls to (approximate) max-flow. Naïvely, these max-flows can be computed recursively, resulting in a congestion approximator algorithm that makes recursive calls to max-flow, similar to [RST14]. Our key insight is that these max-flow instances are actually well-structured enough that recursion is unnecessary. In particular, to construct partition $\mathcal{P}_{i+1}$, the first $i$ partitions $\mathcal{P}_{1},\mathcal{P}_{2},\ldots,\mathcal{P}_{i}$—which the algorithm has already computed—can be converted to a pseudo-congestion approximator, which then translates to a max-flow algorithm sufficient for these well-structured instances.
@@ 3 Preliminaries
We are given an undirected, capacitated graph $G=(V,E)$. The graph has $n$ vertices and $m$ edges, and each edge $e\in E$ has capacity $c(e)$ in the range $[1,W]$. For a set $C\subseteq V$, let $\partial C$ denote the set of edges with exactly one endpoint in $C$, and let $\delta C=\sum_{e\in\partial C}c(e)$ denote the total capacity of edges in $\partial C$. For a vertex $v\in V$, let $\deg(v)$ denote the capacitated degree of vertex $v$, which is also equal to $\delta\{v\}$. We sometimes write $\partial_{G}C$, $\delta_{G}C$, and $\deg_{G}(v)$ to emphasize that the values are with respect to graph $G$.
@@
Given a graph $H=(U,F)$, a congestion-approximator $\mathcal{C}$ of quality $\alpha$ is a collection of subsets of $U$ such that for any demand $\mathbf{b}$ satisfying $|\mathbf{b}(C)|\leq\delta_{H}C$ for all $C\in\mathcal{C}$, there is a flow routing demand $\mathbf{b}$ with congestion $\alpha$. Through Sherman’s framework [She13, She17], a congestion-approximator of quality $\alpha$ translates to a $(1+\epsilon)$-approximate maximum flow algorithm with running time $\tilde{O}(\epsilon^{-1}\alpha m)$. [^12] The $\tilde{O}(\cdot)$ notation hides polylogarithmic factors in $n$. In our paper, it is most convenient to work with a stronger variant of approximate min-cut/max-flow called fair cut/flow [LNPS23], which is formally defined in Section 5.2.
@@ 4 Bottom-Up Congestion-Approximator
In this section, we formally state our congestion-approximator construction and prove its quality guarantee. See Figure 1 for an illustration. The most important distinction is that we define and analyze mixing on vertex weightings, not edges. This simplifies the notation since we can avoid working with the subdivision graph in [RST14]. For example, in property (2) below, the mixing of the vertex weighting $\mathrm{deg}_{\partial\mathcal{P}_{i}\cup\partial C}|_{C}$ is conceptually equivalent to the mixing of the edges of $\partial\mathcal{P}_{i}$ internal to $C$ together with the boundary edges $\partial C$.
@@
Then, their union $\mathcal{C}=\bigcup_{i\in[L]}\mathcal{R}_{\geq i}$ is a congestion-approximator with quality $16L^{2}\alpha\beta$.
@@
Consider their union $\mathcal{C}=\bigcup_{i\in[L]}\mathcal{R}_{\geq i}$. Then, for any demand $\mathbf{b}\in\mathbb{R}^{V}$ satisfying $|\mathbf{b}(C)|\leq\delta C$ for all $C\in\mathcal{C}$, there exists a demand $\mathbf{b}^{\prime}\in\mathbb{R}^{V}$ satisfying $|\mathbf{b}^{\prime}|\leq\deg_{\partial\mathcal{P}_{L}}$ and a flow routing demand $\mathbf{b}-\mathbf{b}^{\prime}$ with congestion $16L^{2}\alpha\beta$.
@@
Consider partitions $\mathcal{P}_{1},\mathcal{P}_{2},\ldots,\mathcal{P}_{L}$ that satisfy the assumptions of Theorem 4.1. For a given demand $\mathbf{b}\in\mathbb{R}^{V}$ satisfying $|\mathbf{b}(C)|\leq\delta C$ for all $C\in\mathcal{C}$, we want to establish a flow routing demand $\mathbf{b}$ with congestion $16L^{2}\alpha\beta$. Theorem 4.1 then follows from the definition of congestion-approximator.
Apply Lemma 4.2 to the partitions $\mathcal{P}_{1},\mathcal{P}_{2},\ldots,\mathcal{P}_{L}$ and the demand $\mathbf{b}$. We obtain a demand $\mathbf{b}^{\prime}\in\mathbb{R}^{V}$ satisfying $|\mathbf{b}^{\prime}|\leq\deg_{\partial\mathcal{P}_{L}}$ and a flow $f$ routing demand $\mathbf{b}-\mathbf{b}^{\prime}$ with congestion $16L^{2}\alpha\beta$. By assumption (1) of Theorem 4.1, we have $\mathcal{P}_{L}=\{V\}$, which implies that $\partial\mathcal{P}_{L}=\emptyset$. Since $|\mathbf{b}^{\prime}|\leq\deg_{\partial\mathcal{P}_{L}}=\mathbf{0}$, we must have $\mathbf{b}^{\prime}=\mathbf{0}$. It follows that the flow $f$ routes demand $\mathbf{b}$ with congestion $16L^{2}\alpha\beta$, finishing the proof. ∎
@@
Let $\mathbf{b}\in\mathbb{R}^{V}$ be a flow demand satisfying $|\mathbf{b}(C)|\leq\delta C$ for all $C\in\mathcal{C}$. We need to construct a demand $\mathbf{b}^{\prime}\in\mathbb{R}^{V}$ satisfying $|\mathbf{b}^{\prime}|\leq\deg_{\partial\mathcal{P}_{L}}$ and a flow routing demand $\mathbf{b}-\mathbf{b}^{\prime}$ with congestion $16L^{2}\alpha\beta$.
@@
    The flow $f_{i}$ has congestion $16L\alpha\beta$.
@@
    The demand $\mathbf{b}_{i}$ satisfies $|\mathbf{b}_{i}|\leq\deg_{\partial\mathcal{R}_{\geq i+1}}$.
The lemma below shows that properties (1), (2), and (4) alone are sufficient to prove Lemma 4.2 with demand $\mathbf{b}^{\prime}=\mathbf{b}_{L-1}$ and flow $f_{1}+f_{2}+\cdots+f_{L-1}$.
@@
Suppose that properties (1), (2), and (4) hold for each $i\in[L-1]$. Then, the demand $\mathbf{b}_{L-1}$ satisfies $|\mathbf{b}_{L-1}|\leq\deg_{\partial\mathcal{P}_{L}}$, and the flow $f_{1}+f_{2}+\cdots+f_{L-1}$ routes demand $\mathbf{b}-\mathbf{b}_{L-1}$ with congestion $16L^{2}\alpha\beta$.
@@
Observe that $\partial\mathcal{R}_{\geq L}=\partial\mathcal{P}_{L}$ by definition. The demand $\mathbf{b}_{L-1}$ satisfies $|\mathbf{b}_{L-1}|\leq\deg_{\partial\mathcal{R}_{\geq L}}=\deg_{\partial\mathcal{P}_{L}}$ by property (4) on iteration $i=L-1$. By property (1) over all $i\in[L-1]$, the flow $f_{1}+f_{2}+\cdots+f_{L-1}$ routes demand $(\mathbf{b}_{0}-\mathbf{b}_{1})+(\mathbf{b}_{1}-\mathbf{b}_{2})+\cdots+(\mathbf{b}_{L-2}-\mathbf{b}_{L-1})=\mathbf{b}-\mathbf{b}_{L-1}$. The congestion is at most the sum of the congestions of each $f_{i}$, which is $(L-1)\cdot 16L\alpha\beta\leq 16L^{2}\alpha\beta$ by property (2). ∎
@@
    $|\mathbf{s}|\leq\deg_{\partial\mathcal{R}_{\geq i}}$.
@@
    The flow $f$ has congestion $16L\alpha\beta$.
@@
For the inductive step, we apply Lemma 4.6 on iteration $i\geq 1$ and the vector $\mathbf{s}=\mathbf{b}_{i-1}$. We first verify the conditions on $\mathbf{s}$ required by Lemma 4.6.
@@
    Condition a follows from property (4) for iteration $i-1$, which is assumed inductively.
@@
    Since $\mathbf{s}=\mathbf{b}_{i-1}$, we have $|\mathbf{s}(C)|=|\mathbf{b}_{i-1}(C)|=|\mathbf{b}_{0}(C)|=|\mathbf{b}(C)|$. Finally, since the initial flow demand $\mathbf{b}\in\mathbb{R}^{V}$ satisfies $|\mathbf{b}(C)|\leq\delta C$, we have $|\mathbf{s}(C)|=|\mathbf{b}(C)|\leq\delta C$, establishing condition b.
With the conditions fulfilled, Lemma 4.6 outputs a flow $f$ which we set as $f_{i}$, immediately satisfying property (2). We set $\mathbf{b}_{i}=\mathbf{t}$ so that flow $f_{i}$ routes demand $\mathbf{s}-\mathbf{t}=\mathbf{b}_{i-1}-\mathbf{b}_{i}$ and property (1) holds. Properties (3) and (4) follow from properties iii and i of Lemma 4.6, respectively, completing the induction and hence the proof. ∎
For the rest of this section, we establish Lemma 4.6, the most technical part of the proof. We first establish a helper claim about constructing certain demands and flows. The proof is quite technical and is split into several subclaims (and their proofs).
@@
For any $i\in[L-1]$, consider any vector $\mathbf{x}\in\mathbb{R}^{V}$ with $|\mathbf{x}|\leq\deg_{\partial\mathcal{R}_{\geq i}}$. There exists a vector $\mathbf{y}\in\mathbb{R}^{V}$ such that
  1. 1.
    $|\mathbf{y}|\leq 2\deg_{\partial\mathcal{P}_{i}}+2L\beta\deg_{\partial\mathcal{P}_{i+1}}$.
  2. 2.
  3. 3.
    There is a flow routing demand $\mathbf{x}-\mathbf{y}$ with congestion $4L\beta$.
Proof.
We begin with a few subclaims.
Subclaim 4.9.
@@
Subclaim 4.10.
For any $i\in[L-1]$, consider any vector $\mathbf{x}\in\mathbb{R}^{V}_{\geq 0}$ with $\mathbf{x}\leq\deg_{\partial\mathcal{R}_{\geq i}}$. There exists a vector $\mathbf{y}\in\mathbb{R}^{V}_{\geq 0}$ with $\mathbf{y}\leq 2\deg_{\partial\mathcal{P}_{i}}$ and a flow routing demand $\mathbf{x}-\mathbf{y}$ with congestion $(2L-2i)\beta$.
@@
We prove the statement by induction from $i=L$ down to $i=1$. For the base case $i=L$, since $\mathcal{R}_{\geq L}=\mathcal{P}_{L}$, we can simply set $\mathbf{y}=\mathbf{x}$ and take the empty flow.
@@
For the inductive step, consider a vector $\mathbf{x}\in\mathbb{R}^{V}_{\geq 0}$ with $\mathbf{x}\leq\deg_{\partial\mathcal{R}_{\geq i}}$. Define $\mathbf{x}^{\prime}\in\mathbb{R}^{V}_{\geq 0}$ as
@@
$\mathbf{x}^{\prime}(v)=\begin{cases}\displaystyle\frac{\deg_{\partial\mathcal{R}_{\geq i+1}}(v)}{\deg_{\partial\mathcal{R}_{\geq i}}(v)}\cdot\mathbf{x}(v)&\text{if }\deg_{\partial\mathcal{R}_{\geq i}}(v)>0,\\ 0&\text{otherwise},\end{cases}$